A fuel pump sends gasoline from a car's fuel tank
Answer #1

[volume] = [mass] / [density] so the volume flow rate of the gasoline is just [mass flow rate] / density

= [5.79 * 10^-2 kg/s] / [735 kg/m^3]
=7.877 * 10^-5 m^3 /s

The cross sectional area (A) of the fuel line of radius r is given by
A = πr² = π * (2.95 * 10^-3)²
A = 2.73 * 10^-5 m²

Linear flow rate = [volume flow rate] / [cross sectional area]
= [7.877 * 10^-5 m^3/s] / [2.73 * 10^-5 m^2]
= 2.88 m/s

Calculation may be wrong but concept is right